# Field-based Molecule Generation

### v1

@@ Appendix I Explicit aromatic bonds and neutrality
Figure 14: Three aromatic molecules, with increasingly harder to determine neutrality. In the first case
@@
We have used the molecules reported in the MiDi github repository. We note that on the QM9 dataset, both our method and MiDi generate molecules in Kekulé form, and they are directly comparable in Table 1. For the GEOM neutrality results in Table 3, we have reported an upper-bound for MiDi, by considering bond type “aromatic” as 1.5 electro occupancy and allowing C to have both 0 and -0.5 formal charge. Since many N atoms in GEOM are (in actuality) ionized, it is not ”allowed” to have a formal charge of -1. Regardless, the comparison is very hard to do, and it is difficult to say which model generates more neutral atoms or molecules.
@@
We compute the maximum test similarity using valid molecules. This results in a bias in the methods that generate aromatic bonds explicitly, as RDKit will require these molecular graphs to be converted to Kekulé representations. In this case, this metric will not consider molecules containing incorrect aromatic rings (as determined by RDKit’s SanitizeMol). However, in our case, since we directly generate Kekulé molecule representations, this quality filter does not exist anymore, and samples that have incorrect aromatic rings (but are not identified as aromatic because of e.g., one circle atom being incorrectly hybridized) will be considered in the final metric. Note how this bias is unrelated to the performance of the generative model, as those unkekulizable molecules are still generated by explicit-aromatic generating methods, but are simply filtered out by this metric.

### v2

@@ Appendix I Explicit aromatic bonds and neutrality
Figure 19: Three aromatic molecules, with increasingly harder to determine neutrality.
@@
We have used the molecules reported in the MiDi github repository. We note that on the QM9 dataset, both our method and MiDi generate molecules in Kekulé form, and they are directly comparable in Table 1 (QM9 columns). For the GEOM-Drugs neutrality results in Table 1 (GEOM-Drugs columns), we have reported an upper-bound for MiDi, by considering bond type “aromatic” as 1.5 electron occupancy and allowing C to have both 0 and -0.5 formal charge. Since many N atoms in GEOM-Drugs are (in actuality) ionized, it is not ”allowed” to have a formal charge of -1. Regardless, the comparison is very hard to do, and it is difficult to say which model generates more neutral atoms or molecules.
@@
We compute the maximum test similarity using valid molecules. This results in a bias in the methods that generate aromatic bonds explicitly, as RDKit will require these molecular graphs to be converted to Kekulé representations (Table 2). In this case, this metric will not consider molecules containing incorrect aromatic rings (as determined by RDKit’s SanitizeMol). However, in our case, since we directly generate Kekulé molecule representations, this quality filter does not exist anymore, and samples that have incorrect aromatic rings (but are not identified as aromatic because of e.g., one circle atom being incorrectly hybridized) will be considered in the final metric. Note how this bias is unrelated to the performance of the generative model, as those unkekulizable molecules are still generated by explicit-aromatic generating methods, but are simply filtered out by this metric.
Appendix J Complexity analysis
J.0.1 Time scaling
We use convolutional layers as the main blocks of our UNet parametrization. As such, the complexity of our NN scales linearly with $H\cdot W\cdot D$. We omit the channel term $C_{\text{in}}\cdot C_{\text{out}}$, as it is equivalent to the hidden activations on node states from point-cloud methods. Linear attention layers are added after each UNet layer, giving $\mathcal{O}(H\cdot W\cdot D)$.
Comparatively, all layers of point-cloud DDPM methods scale quadratically in the number of atoms $\mathcal{O}(N^{2})$. The scaling of these two methods is not directly comparable, since the number of atoms does not directly relate to the necessary volume that encompasses all molecules of that size, which is highly dataset-dependent. Still, a theoretical upper limit can be determined.
Consider a dataset that contains molecules of atoms, which, for simplicity, are all C atoms bonded through covalent bond type 1 (approximately Å). We now discuss how the atoms should be arranged to create the worst possible scenario (without considering physical constraints). The upper limit of our grid sizes would be achieved when all three of the following n-atom molecules are present:
  • One molecule has all atoms placed linearly, which will determine $H=N\cdot 1.35$Å, and $W=D=0$.
  • One molecule is perfectly planar, having $H=W=\frac{N\cdot 1.35}{2}$ and D=0. Note that $W>H$ is not possible, since our alignment method (Section 3.5) would rotate the molecule s.t. the longest axis is $H$.
  • One molecule has $H=W=D=\frac{N\cdot 1.35}{3}$
The scaling of our convolutional layers would be $\mathcal{O}(\frac{1.35N}{r}\cdot\frac{1.35N}{2r}\cdot\frac{1.35N}{3r}$), where $r$ is the resolution of our evaluated field locations. We compare the number of floating point operations in terms of the number of atoms in Figure 20. For FMG, we show both the theoretical maximum and the grids that encompass GEOM-Drugs molecules of various numbers of atoms. In Figure 21, we determine the training and generation time (in seconds) of FMG and EDM.
When close to the theoretical upper limit, our method becomes intractable, but this case is highly unrealistic (Figure 20, comparing ”FMG max” with ”FMG GEOM”). In Figure 2 a), FMG has a comparable generation time but takes longer to train. Figure 2 b) depicts our training time for various molecule sizes, where 99.7% of the molecules are kept (used in this work). The molecule graphs were chosen s.t. they fit within a grid size of $64\cdot 40\cdot 32$.
Figure 20: Number of floating point operations for EDM and FMG over the number of atoms in a molecule. The theoretical maximum of FMG differs substantially from the real datasets, and its practicality relies on application specifications. For the axis lengths in FMG 0.33Å was used (same value as the one used in our model).
Figure 21: Time elapsed (in seconds) for EDM and FMG models, over various molecule sizes (in number of atoms, and tensor field volumes, for EDM and FMG, respectively). We used 100 iterations with a batch of four molecules to determine the training curves. For this experiment, we used an Nvidia A100 GPU. a) All GEOM-Drugs molecules are considered. b) Only molecules used in the current work (99.7% of the molecules) that fit within our grid size of 64·40·32 are used in the analysis.
J.0.2 Memory Scaling
Increasing the molecule sizes has a significant impact on memory overhead. Similar to the analysis done in the time scaling analysis in Appendix J.0.1, we construct grids that can fit all GEOM-Drugs molecules with a certain number of atoms $N$. We report the maximum memory allocated for EDM (Hoogeboom et al., 2022) and our method, when generating samples of various atom sizes (determined on the x label by the number of atoms $N$, or dots denoting volume sizes), with 32 samples per batch.
Figure 22 a) shows that our method performs better for larger molecules, but worse for smaller ones. We note here that GNN architecture batching is done by considering a single large graph, where nodes from different batches are disconnected. Therefore, adjacency matrices of this implementation have poor scaling for large batch sizes: $((\sum_{b\in B}{N_{b}})^{2})$, where $B$ is the batch sizes and $N_{b}$ is the number of atoms in a molecule.
We find that increasing the number of unique channels (atoms or bonds) is negligible and linear. The additional memory required stems from the input CNN layer’s kernel dimension (the depth sizes of the initial 3D CNN kernels are equal to $|\mathcal{A}|+|\mathcal{B}|$), and the larger intermediate states. However, these are considerably smaller than the architecture’s overhead (Figure 22 b). Channel scaling is performed with a batch size of 1.
Figure 22: Memory scaling in terms of molecule size for generating molecules. a) We report the maximum GPU memory allocated when sampling for various grid sizes, which fit all molecules of the corresponding number of atoms. We additionally report the memory when using the grid sizes chosen for this work. b) Generation memory in terms of the number of channels.